# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ Abstract
Infinite-dimensional, holomorphic functions have been studied in detail over the last several decades, due to their relevance to parametric differential equations and computational uncertainty quantification. The approximation of such functions from finitely many samples is of particular interest, due to the practical importance of constructing surrogate models to complex mathematical models of physical processes. In a previous work, [6] we studied the approximation of so-called Banach-valued, $(\bm{b},\varepsilon)$-holomorphic functions on the infinite-dimensional hypercube $[-1,1]^{\mathbb{N}}$ from $m$ (potentially adaptive) samples. In particular, we derived lower bounds for the adaptive $m$-widths for classes of such functions, which showed that certain algebraic rates of the form $m^{1/2-1/p}$ are the best possible regardless of the sampling-recovery pair. In this work, we continue this investigation by focusing on the practical case where the samples are pointwise evaluations drawn identically and independently from the underlying probability measure for the problem. Specifically, for Hilbert-valued $(\bm{b},\varepsilon)$-holomorphic functions, we show that the same rates can be achieved (up to a small polylogarithmic or algebraic factor) for essentially arbitrary tensor-product Jacobi measures. Our reconstruction maps are based on least squares and compressed sensing procedures using the corresponding orthonormal Jacobi polynomials. In doing so, we strengthen and generalize past work that has derived weaker nonuniform guarantees for the uniform and Chebyshev measures (and corresponding polynomials) only. We also extend various best $s$-term polynomial approximation error bounds to arbitrary Jacobi polynomial expansions. Overall, we demonstrate that i.i.d. pointwise samples drawn from an underlying probability measure are near-optimal for the recovery of infinite-dimensional, holomorphic functions.
Keywords: high-dimensional approximation, holomorphic functions, pointwise samples, information complexity, Hilbert spaces
@@ 1 Introduction
In many applications in biology, chemistry, economics, engineering and elsewhere, a mathematical model is represented as the solution of a (system of) parametric Differential Equations (DEs), where the input variable $\bm{y}\in\mathcal{U}$ belongs to a subset of $\mathbb{R}^{\mathbb{N}}$ instead of a finite-dimensional space and the output is an element of an infinite-dimensional function space, e.g., a Hilbert or Banach space $\mathcal{V}$. Notably, this is the case for certain coefficients in parametric DE that are infinite-dimensional random fields represented by a Karhunen–Loève expansion [9, 14, 34, 20].
@@
In many applications, working directly with such a model is computationally infeasible, since each model evaluation requires a costly DE solve [37, 18]. Therefore, a standard approach involves constructing a surrogate model, i.e., an approximation $\hat{f}$ with certain error guarantees measuring, in a suitable norm, how far the approximation is to the true target function $f:\bm{y}\in\mathcal{U}\mapsto f(\bm{y})\in\mathcal{V}$. Since it is often undesirable to alter the existing numerical DE code, one strives to do this via a nonintrusive method, where the approximation $\hat{f}$ is computed from a collection of pointwise samples (also known as snapshots). However, this is a challenging task. The number of samples available to produce such approximations is often highly limited, and since the underlying function is infinite dimensional, the so-called curse of dimensionality is an ever-present issue.
@@
Fortunately, in the context of parametric DEs, many solution maps $\bm{y}\mapsto f(\bm{y})$ are smooth functions. Specifically, a long series of works have shown that solution maps of various different parametric DEs belong to the class of $(\bm{b},\varepsilon)$-holomorphic functions (see [13] and [5, Chpt. 4] for reviews). Here $\bm{b}\in[0,\infty)^{\mathbb{N}}$ is a nonnegative sequence, whose $i$th entry is associated with the smoothness in the $i$th variable, and $\varepsilon>0$ is a scalar. Because of the relevance of this class to surrogate model construction, it is therefore crucial to understand the fundamental limits of how well one can approximate such functions from finitely many samples. In our previous work [6], we derived lower bounds for the (adaptive) $m$-widths for classes of $(\bm{b},\varepsilon)$-holomorphic functions. In other words, we examined how well one can approximate such functions using $m$ arbitrary (adaptive) linear measurements and arbitrary reconstruction maps. Our results considered the case where $\bm{b}$ (or its monotone majorant) is $\ell^{p}$-summable for some $10<p<$ 1, this being a standard assumption in the above literature. In particular, we showed that, when the $m$-width decays, it does so with rate no faster than the algebraic power $m^{1/2-1/p}$.
@@
Unfortunately, [6] does not directly apply to the surrogate model construction problem, where the measurements are pointwise samples of the target function. The goal of this paper is to continue the investigations of [6] by focusing on the case of pointwise samples (also known as standard information [29, 30]). Specifically, we show that these lower bounds can be achieved (up to either logarithmic or arbitrarily small algebraic factors) from i.i.d. pointwise samples (also known as random information) drawn from arbitrary Jacobi measures. In other words, random information constitutes nearly optimal information for recovering $(\bm{b},\varepsilon)$-holomorphic functions. We also derive explicit reconstruction procedures for attaining these rates, thus demonstrating practical schemes for surrogate model construction from pointwise samples.
@@ 1.1 Contributions and related work
In this work, we study the approximation of $(\bm{b},\varepsilon)$-holomorphic functions $f:\mathcal{U}\rightarrow\mathcal{V}$, where $\mathcal{U}=[-1,1]^{\mathbb{N}}$ and $\mathcal{V}$ is a Hilbert space. See §2.2 for the full definition. Here $\bm{b}\in[0,\infty)^{\mathbb{N}}$ and $\varepsilon>0$ is a (technically redundant) parameter that affects the size of the region of holomorphy. As discussed in [3, 5, 6], $\bm{b}$ controls the anisotropy of such a function. The size of the entries of $\bm{b}$ relate to the smoothness of $f$ with respect to each of its variables. Following our previous work [6], we consider two scenarios. First, the known anisotropy (known $\bm{b}$) case, where the reconstruction map is allowed to depend on $\bm{b}$. Second, the more challenging unknown anisotropy (unknown $\bm{b}$) case, where the reconstruction map is required to be independent of $\bm{b}$. For samples, we focus on pointwise samples that are drawn i.i.d. from tensor-product Jacobi probability measures. Besides a mild boundedness assumption (see Assumption 4.1), these can be arbitrary.
@@
Recently, various works have studied the approximation of $(\bm{b},\varepsilon)$-holomorphic functions from pointwise samples, with similar algebraic rates being established for polynomial-based methods or, recently, deep neural networks [3, 4, 2, 5, 21, 15, 22]. In particular, the book [5] describes approximation of such functions from pointwise samples in the scalar-valued case via Chebyshev and Legendre polynomials using least-squares and compressed sensing techniques. The work [4] extended this to the case of Hilbert-valued functions and showed that the resulting compressed sensing methods can be efficiently implemented in subexponential computational cost. In [2, 3], by emulating polynomial-based least squares and compressed sensing techniques as deep neural network training procedures, it was shown that the same rates approximation rates can be achieved using deep learning. Moreover, [3] also extended these previous results from Hilbert- to Banach-valued functions. The main distinguishing features in this work are:
@@
    We establish uniform upper bounds, i.e., they hold for any function belonging to the given function class.
@@
More precisely, in the known anisotropy case, we combine three main arguments. First, standard least-squares analysis [12]. Second, new bounds for certain weighted best approximation errors for Jacobi polynomial expansions. Third, a particular nested index set construction based on ideas of [25, 26] to achieve the desired rate. Conversely, in the unknown anisotropy case, we construct polynomial approximations via compressed sensing and weighted $\ell^{1}$-minimization. Our technical arguments are based on [4]. However, we modify these to derive stronger uniform error guarantees, and also extend them from the case of tensor-product Chebyshev and Legendre polynomials to arbitrary tensor-product Jacobi polynomials and i.i.d. sampling from the corresponding Jacobi measures. While Chebyshev and Legendre polynomials are, arguably, the most commonly used in practice, the generalization to Jacobi polynomials is not only mathematically interesting, it also has practical relevance to computational uncertainty quantification where the parameters follow beta distributions [36].
@@ 2 Preliminaries
This section introduces further notation, definitions, and other preliminaries that are needed later in the paper.
@@ 2.1 Notation
Let $\mathbb{N}$ and $\mathbb{N}_{0}$ be the set of positive and nonnegative integers, respectively. We write $[N]=\{1,2,\ldots,N\}$ for $N\in\mathbb{N}\cup\{\infty\}$, with the convention that $[N]=\mathbb{N}$ if $N=\infty$. Let $\mathbb{R}^{N}$ the vector space of real vectors of length $N\in\mathbb{N}$ and $\mathbb{R}^{\mathbb{N}}$ be the vector space of real sequences indexed over $\mathbb{N}$. In either space, we write $\bm{e}_{j}=(\delta_{j,k})_{k\in[N]}$ for the standard basis vectors, where $j\in\mathbb{N}$ or $j\in[N]$.
For $N\in\mathbb{N}\cup\{\infty\}$, we write $\bm{0}\in\mathbb{N}^{N}_{0}$ and $\bm{1}\in\mathbb{N}^{N}_{0}$ for the multi-indices consisting of all zeros and all ones, respectively. For any multi-indices $\bm{\nu}=(\nu_{i})_{i\in[N]}$ and $\bm{\mu}=(\mu_{i})_{i\in[N]}$, the inequality $\bm{\nu}\leq\bm{\mu}$ (or $\bm{\nu}<\bm{\mu}$) is understood componentwise, i.e., $\bm{\nu}\leq\bm{\mu}$ means that ${\nu_{j}}\leq{\mu_{j}}$ for all $j\in[N]$. We also write
@@
For $1\leq p\leq\infty$ $(0<p<1)$, we write ${\left\|\cdot\right\|}_{p}$ for the usual vector $\ell^{p}$-norm (quasi-norm) on $\mathbb{R}^{N}$, for $N\in\mathbb{N}\cup\{\infty\}$, i.e.,
${\|\bm{c}\|}_{p}=\begin{cases}\left(\sum_{i\in[N]}|{c_{i}}|^{p}\right)^{\frac{1}{p}}&0<p<\infty,\\ \sup_{i\in[N]}|{c_{i}}|&p=\infty,\end{cases}\qquad\forall\bm{c}=(c_{i})_{i\in[N]}\in\mathbb{R}^{N}.$
We write $\ell^{p}([N])$ for the space of sequences for which $\|\bm{c}\|_{p}<\infty$ (see also §5.2 for a more general definition). Given a set $S\subseteq[N]$ and $\bm{c}=(c_{i})_{i\in[N]}\in\mathbb{R}^{N}$, we write $\bm{c}_{S}$ for the vector with $i$th entry equal to $c_{i}$ if $i\in S$ and zero otherwise. We also write $S^{c}$ for the set complement $[N]\setminus S$ of $S$.
Throughout this paper we consider variables $\bm{y}\in\mathcal{U}$, where $\mathcal{U}=[-1,1]^{\mathbb{N}}$. We let $\varrho$ be a probability measure on $\mathcal{U}$ that is obtained as a tensor-product of probability measures on $[-1,1]$, i.e.,
$\varrho=\varrho_{1}\times\varrho_{2}\times\cdots.$ (2.1)
The existence of this measure is guaranteed by the Kolmogorov extension theorem [23, Thm. 14.36]. We also consider a Hilbert space $(\mathcal{V},\langle\cdot,\cdot\rangle_{\mathcal{V}})$ with induced norm ${\left\|\cdot\right\|}_{\mathcal{V}}$.
For $1\leq p\leq\infty$, we define the weighted Lebesgue-Bochner space $L^{p}_{\varrho}(\mathcal{U};\mathcal{V})$ as the space consisting of (equivalence classes of) strongly $\varrho$-measurable functions $f:\mathcal{U}\rightarrow\mathcal{V}$ for which $\|f\|_{L^{p}_{\varrho}(\mathcal{U};\mathcal{V})}<\infty$, where
@@
Throughout this work, we write $C(\cdot)$ for a constant that only depends on its arguments, and which is allowed to vary from one line to the next. On occasion, we write $\widetilde{C}$ or $C_{1}$, $C_{2}$ and so forth to distinguish specific constants. We also write lower-case $c$ for a universal constant (i.e., a constant that is independent of any variables).
@@ 2.2 The space of $(\boldsymbol{b},\varepsilon)$-holomorphic functions
$\mathcal{H}(\bm{b})=\left\{f:\mathcal{U}\rightarrow\mathcal{V}\text{ $(\bm{b},1)$-holomorphic}:{\|f\|}_{L^{\infty}(\mathcal{R}(\bm{b});\mathcal{V})}{=:\operatorname*{ess\,sup}_{\bm{z}\in\mathcal{R}(\bm{b})}{\|f(\bm{z})\|}_{\mathcal{V}}}\leq 1\right\}.$ (2.5)
For a further discussion on $(\bm{b},\varepsilon)$-holomorphic functions, see [5, 11, 35].
@@
$\mathcal{H}(p)=\bigcup\left\{\mathcal{H}(\bm{b}):\bm{b}\in\ell^{p}(\mathbb{N}),\ \bm{b}\in[0,\infty)^{\mathbb{N}},\ {\|\bm{b}\|}_{p}\leq 1\right\},$

### v3

@@ Abstract
Infinite-dimensional, holomorphic functions have been studied in detail over the last several decades, due to their relevance to parametric differential equations and computational uncertainty quantification. The approximation of such functions from finitely-many samples is of particular interest, due to the practical importance of constructing surrogate models to complex mathematical models of physical processes. In a previous work, [6] we studied the approximation of so-called Banach-valued, $(\bm{b},\varepsilon)$-holomorphic functions on the infinite-dimensional hypercube $[-1,1]^{\mathbb{N}}$ from $m$ (potentially adaptive) samples. In particular, we derived lower bounds for the adaptive $m$-widths for classes of such functions, which showed that certain algebraic rates of the form $m^{1/2-1/p}$ are the best possible regardless of the sampling-recovery pair. In this work, we continue this investigation by focusing on the practical case where the samples are pointwise evaluations drawn identically and independently from the underlying probability measure for the problem. Specifically, for Hilbert-valued $(\bm{b},\varepsilon)$-holomorphic functions, we show that the same rates can be achieved (up to a small polylogarithmic or algebraic factor) for tensor-product Jacobi measures. Our reconstruction maps are based on least squares and compressed sensing procedures using the corresponding orthonormal Jacobi polynomials. In doing so, we strengthen and generalize past work that has derived weaker nonuniform guarantees for the uniform and Chebyshev measures (and corresponding polynomials) only. We also extend various best $s$-term polynomial approximation error bounds to arbitrary Jacobi polynomial expansions. Overall, we demonstrate that i.i.d. pointwise samples drawn from an underlying probability measure are near-optimal for the recovery of infinite-dimensional, holomorphic functions.
Keywords: high-dimensional approximation, holomorphic functions, pointwise samples, information complexity
@@ 1 Introduction
In many applications in biology, chemistry, economics, engineering and elsewhere, a mathematical model is represented as the solution of a (system of) parametric Differential Equations (DEs), where the input variable $\bm{y}\in\mathcal{U}$ belongs to a subset of $\mathbb{R}^{\mathbb{N}}$ instead of a finite-dimensional space and the output is an element of an infinite-dimensional function space, e.g., a Hilbert or Banach space $\mathcal{V}$. Notably, this is the case in parametric DEs when infinite-dimensional random fields are represented by their Karhunen–Loève expansions [9, 14, 34, 20].
@@
Working directly with such a model is often infeasible, since each model evaluation requires a costly DE solve [37, 18]. Therefore, a standard approach is to construct surrogate model, i.e., an approximation $\hat{f}$ to the true target function $f:\bm{y}\in\mathcal{U}\mapsto f(\bm{y})\in\mathcal{V}$. Since it is often undesirable to alter the existing numerical DE code, one typically strives to do this via a nonintrusive method, where the approximation $\hat{f}$ is computed from a collection of pointwise samples of $f$ (also known as snapshots). However, this is a challenging task. The number of samples available to produce such approximations is often highly limited, and since the underlying function is infinite dimensional, the so-called curse of dimensionality is an ever-present issue.
@@
Fortunately, solution maps $\bm{y}\mapsto f(\bm{y})$ arising from parametric DEs are often smooth functions. A long series of works have shown that solution maps of various different parametric DEs belong to the class of $(\bm{b},\varepsilon)$-holomorphic functions (see [13] and [5, Chpt. 4] for reviews). Here $\bm{b}\in[0,\infty)^{\mathbb{N}}$ is a nonnegative sequence, whose $i$th entry is associated with the smoothness in the $i$th variable, and $\varepsilon>0$ is a scalar. Because of the relevance of this class to surrogate models, it is crucial to understand the fundamental limits of how well one can approximate such functions from finitely-many samples. In our previous work [6], we derived lower bounds for the (adaptive) $m$-widths for classes of $(\bm{b},\varepsilon)$-holomorphic functions. In other words, we examined how well one can approximate such functions using $m$ arbitrary (adaptive) linear measurements and arbitrary reconstruction maps. Our results considered the case where $\bm{b}$ (or its monotone majorant) is $\ell^{p}$-summable for some $10<p<$ 1, this being a standard assumption in the above literature. In particular, we showed that, when the $m$-width decays, it does so with rate no faster than the algebraic power $m^{1/2-1/p}$.
@@
Unfortunately, [6] does not directly apply to the surrogate model construction problem, where the measurements are pointwise samples of the target function. The goal of this paper is to continue the investigations of [6] by focusing on the case of pointwise samples (also known as standard information [29, 30]). We show that these lower bounds can be achieved (up to either logarithmic or arbitrarily small algebraic factors) from i.i.d. pointwise samples (also known as random information) drawn from arbitrary Jacobi measures. In other words, random information constitutes nearly optimal information for recovering $(\bm{b},\varepsilon)$-holomorphic functions. We also derive explicit reconstruction procedures that attain these rates, thus demonstrating practical schemes for surrogate model construction from pointwise samples.
@@ 1.1 Contributions and related work
In this work, we study the approximation of $(\bm{b},\varepsilon)$-holomorphic functions $f:\mathcal{U}\rightarrow\mathcal{V}$, where $\mathcal{U}=[-1,1]^{\mathbb{N}}$ and $\mathcal{V}$ is a Hilbert space. See §2.2 for the full definition. As discussed in [4, 5, 6], $\bm{b}$ controls the anisotropy of such a function. The size of the entries of $\bm{b}$ relate to the smoothness of $f$ with respect to each of its variables. Following our previous work [6], we consider two scenarios. First, the known anisotropy (known $\bm{b}$) case, where the reconstruction map is allowed to depend on $\bm{b}$. Second, the more challenging unknown anisotropy (unknown $\bm{b}$) case, where the reconstruction map is required to be independent of $\bm{b}$. For samples, we focus on pointwise samples that are drawn i.i.d. from tensor-product Jacobi probability measures. Besides a mild boundedness assumption (see Assumption 4.1), these can be arbitrary.
@@
Recently, various works have studied the approximation of $(\bm{b},\varepsilon)$-holomorphic functions from pointwise samples, with similar algebraic rates being established for polynomial-based methods or, recently, deep neural networks [4, 3, 2, 5, 21, 15, 22]. In particular, [5] describes approximation of such functions from pointwise samples in the scalar-valued case via Chebyshev and Legendre polynomials using least-squares and compressed sensing techniques. The work [3] extended this to the case of Hilbert-valued functions and showed that the resulting compressed sensing methods can be efficiently implemented in subexponential time. In [2, 4], by emulating polynomial-based techniques as deep neural network training procedures, it was shown that the same rates approximation rates can be achieved using deep learning. Moreover, [4] also extended these previous results from Hilbert- to Banach-valued functions. The main distinguishing features in this work are:
@@
    We establish uniform upper bounds, i.e., bounds that hold for any function belonging to the given function class.
@@
More precisely, in the known anisotropy case, we combine three main arguments. First, standard least-squares analysis [12]. Second, new bounds for certain weighted best approximation errors for Jacobi polynomial expansions. Third, a particular nested index set construction based on ideas of [25, 26] to achieve the desired rate. In the unknown anisotropy case, we construct polynomial approximations via compressed sensing and weighted $\ell^{1}$-minimization. Our technical arguments are based on [3]. However, we modify these to derive stronger uniform error guarantees, and extend them from tensor-product Chebyshev and Legendre polynomials to arbitrary tensor-product Jacobi polynomials and i.i.d. sampling from the corresponding Jacobi measures. While Chebyshev and Legendre polynomials are, arguably, the most commonly used in practice, the generalization to Jacobi polynomials is not only mathematically interesting, it also has practical relevance to computational uncertainty quantification where the parameters follow beta distributions [36].
@@ 2 Preliminaries
This section introduces further notation, definitions and other preliminaries that are needed later.
@@ 2.1 Notation
Let $\mathbb{N}$ and $\mathbb{N}_{0}$ be the sets of positive and nonnegative integers, respectively. We write $[N]=\{1,2,\ldots,N\}$ for $N\in\mathbb{N}\cup\{\infty\}$, with the convention that $[N]=\mathbb{N}$ if $N=\infty$. Let $\mathbb{R}^{N}$ be the vector space of real vectors of length $N\in\mathbb{N}$ and $\mathbb{R}^{\mathbb{N}}$ be the vector space of real sequences indexed over $\mathbb{N}$. In either space, we write $\bm{e}_{j}=(\delta_{j,k})_{k\in[N]}$ for the standard basis vectors, where $j\in\mathbb{N}$ or $j\in[N]$.
For $N\in\mathbb{N}\cup\{\infty\}$, we write $\bm{0}\in\mathbb{N}^{N}_{0}$ and $\bm{1}\in\mathbb{N}^{N}_{0}$ for the multi-indices consisting of all zeros and all ones, respectively. For any multi-indices $\bm{\nu}=(\nu_{i})_{i\in[N]}$ and $\bm{\mu}=(\mu_{i})_{i\in[N]}$, the inequality $\bm{\nu}\leq\bm{\mu}$ (or $\bm{\nu}<\bm{\mu}$) is understood componentwise, i.e., $\bm{\nu}\leq\bm{\mu}$ if ${\nu_{j}}\leq{\mu_{j}}$ for all $j\in[N]$. We also write
@@
For $1\leq p\leq\infty$ $(0<p<1)$, we write ${\left\|\cdot\right\|}_{p}$ for the usual vector $\ell^{p}$-norm (quasi-norm) on $\mathbb{R}^{\mathbb{N}}$, and $\ell^{p}(\mathbb{N})$ for the space of sequences for which $\|\bm{c}\|_{p}<\infty$ (see §5.3 for a more general definition). Given a set $S\subseteq[N]$ and $\bm{c}=(c_{i})_{i\in[N]}\in\mathbb{R}^{N}$, we write $\bm{c}_{S}$ for the vector with $i$th entry equal to $c_{i}$ if $i\in S$ and zero otherwise. We also write $S^{c}$ for the set complement $[N]\setminus S$ of $S$.
Let $\bm{y}\in\mathcal{U}$ denote the variable in $\mathcal{U}=[-1,1]^{\mathbb{N}}$ and $\varrho$ be a probability measure on $\mathcal{U}$. Let $(\mathcal{V},\langle\cdot,\cdot\rangle_{\mathcal{V}})$ be a Hilbert space with induced norm ${\left\|\cdot\right\|}_{\mathcal{V}}$. For $1\leq p\leq\infty$, we define the weighted Lebesgue-Bochner space $L^{p}_{\varrho}(\mathcal{U};\mathcal{V})$ as the space consisting of (equivalence classes of) strongly $\varrho$-measurable functions $f:\mathcal{U}\rightarrow\mathcal{V}$ for which $\|f\|_{L^{p}_{\varrho}(\mathcal{U};\mathcal{V})}<\infty$, where
@@
Finally, we write $C(\cdot)$ for a constant that only depends on its arguments, and which is allowed to vary from one line to the next. On occasion, we write $\widetilde{C}$ or $C_{1}$, $C_{2}$ and so forth to distinguish specific constants. We also write lower-case $c$ for a universal constant (i.e., a constant that is independent of any variables).
@@ 2.2 The space of $(\boldsymbol{b},\varepsilon)$-holomorphic functions
$\mathcal{H}(\bm{b})=\left\{f:\mathcal{U}\rightarrow\mathcal{V}\text{ $(\bm{b},1)$-holomorphic}:{\|f\|}_{L^{\infty}(\mathcal{R}(\bm{b});\mathcal{V})}:=\operatorname*{ess\,sup}_{\bm{z}\in\mathcal{R}(\bm{b})}{\|f(\bm{z})\|}_{\mathcal{V}}\leq 1\right\}.$ (2.4)
See [5, 11, 35] for further discussion on $(\bm{b},\varepsilon)$-holomorphic functions.
@@
$\mathcal{H}(p)=\bigcup\left\{\mathcal{H}(\bm{b}):\bm{b}\in\ell^{p}(\mathbb{N}),\ \bm{b}\in[0,\infty)^{\mathbb{N}},\ {\|\bm{b}\|}_{p}\leq 1\right\}$
@@
Note that it is a short argument to show that $\mathcal{H}(p,\mathsf{M})$ is, in fact, equivalent to the union over all $\bm{b}\in\ell^{p}(\mathbb{N})$ that are monotonically decreasing.